OpenStudy (anonymous):

integrate 2+sin(2*sqrt(x)) dx

sam (.sam.):

\[\int\limits \left(2+\sin \left(2 \sqrt{x}\right)\right) \, dx\] Let u=\(\sqrt{x}\) \[du=\frac{1}{2\sqrt{x}}dx\] \[2\sqrt{x}du=dx\] \[\int\limits 2+\sin(2u)[2\sqrt{x}du] \\ \\ \int\limits 2+2usin(2u)du \\ \\ 2 \int\limits 1du+2 \int\limits usin(2u)du\] Then integration by parts for the right integral

OpenStudy (anonymous):

thanks SAM alot :)